When the Cu2+ concentration is 5.57×10-4 M, the observed cell potential at 298K for an electrochemical cell with the following reaction is 1.419V. What is the Mn2+ concentration? Cu2+(aq) + Mn(s) Cu(s) + Mn2+(aq)

Answer: _____M
Answer #1

Cu 2+ (aq) + Mu(s) as cu()+ MM2+ (aq) Ecell = Ecathode - Eanode = Ecurt Icu - Emmst/Mn = 0.34 - 67.18) Ecell = +1.52ul Ave de-0.101 :-0.02955 = log [Mn 2+] (5.57x104) 3.418 - Loq [Mu2t] 9.15.54 x 10-4) anti log (3-418) - [numz+] S.57 x1004 2618-18 X